An experienced accountant can balance the books twice as fast as a new accountant. Working together it takes the accountants 6 hours. How long would it take the experienced accountant working alone?
Answer by checkley75(3666)

x*2x/x+2x=6
2x^2/3x=6
2x^2=18x
2x^2-18x=0
2x(x-9)=0
x-9=0
x=9 hours for the fast accountant.
2*9=18 hours for the slow accountant.
proof
2*9^2/3*9=6
2*81/27=6
162/27=6
6=6